OpenStudy (anonymous):

thanks for the help, its needed! the angle of elevation to the top a of skyscraper is measured to be at 2 degrees from the point on the ground one mile from the building. How tall is the skyscraper?

OpenStudy (goformit100):

▬Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.▬

OpenStudy (imstuck):

You don't need the picture if you draw it out yourself, goformit100. You would use the cosine of 2 degrees. It's elementary, really.

OpenStudy (imstuck):



OpenStudy (imstuck):

Are you sure it's 2 degrees?

OpenStudy (anonymous):

positive that its 2 degrees, thank you for the help

OpenStudy (imstuck):

Ok, then, let's get to work. Are you familiar with the cosine function of an angle?

OpenStudy (anonymous):

yes, im familiar with the cosine function of an angle

OpenStudy (imstuck):

Do you have a calculator? And secondly, notice that I changed miles into feet because it is more normal to measure the height of a building in feet as opposed to miles. Calculator?

OpenStudy (anonymous):

yes, i have a calculator ready at hand! And i've noticed, i believed that would be a lot easier and normal as feet

OpenStudy (imstuck):

Make sure, please, that your calculator is in degrees as opposed to radians! Ok?

OpenStudy (anonymous):

yup, got it

OpenStudy (imstuck):

Cos 2 = side opposite / side adjacent, right?

OpenStudy (imstuck):

so our set up will look like this:\[\cos2=\frac{ x }{ 5280 }\]What we are solving for is x, which is the height of the building. Can you solve that for x?

OpenStudy (imstuck):

Are you still with me?

OpenStudy (anonymous):

yes, i am. Just a little over thinking, is it 2640 ft for the total height of the building?

OpenStudy (imstuck):

No, it's not. Here, let's do this:

OpenStudy (imstuck):

\[\cos(2)=\frac{ x }{ 5280 }\]\[\cos(2) * 5280 = x\]\[\cos(2)=.9993908\]so\[.9993908 * 5280 = 5276.78 ft\]

OpenStudy (imstuck):

See that?

OpenStudy (anonymous):

yes, im following

OpenStudy (imstuck):

So that's your answer...the building is 5276.78 ft tall.

OpenStudy (anonymous):

there is a reason im on this site, ha! And thank you, i can see how easy that really was, I have a bit more confidence in this now

OpenStudy (imstuck):

Very cool!!!